# CroCoDai: A Stablecoin for Cross-Chain Commerce

### v1

@@ Appendix E DSS Parametrization
Table 8: Information about the various collateral types used by the DSS (as of Ethereum block 17121982).
@@
Standard ERC-20 Tokens
ETH-A Ether 233580929 421276897 180% 15000000000 1.50% 145% 4.69%
@@
LP Tokens
UNIV2USDCETH-A USDC/ETH 2139902 0 0% 1.50% 120% 0.04%
@@
Real-World Assets
RWA001-A 6s Capital 14348036 15000000 105% 3.00% 0.29%
@@
PSMs
PSM-USDC-A USD Coin 2118503205 2608700392 123% 10000000000 42.58%
@@
Appendix F Price Data Table
@@
Table 9: Data periods (times are in Greenwich Mean Time).
@@
Figure 6: The correlations between all 100 tokens in our dataset.
Appendix G Portfolio Compositions
@@
Table 10: DSS-based portfolio compositions.
@@
Table 11: Optimal portfolio compositions.
@@
Table 12: Optimal simulation portfolio if other stablecoins are included.

### v4

@@ Appendix B DSS Parametrization
Table 5. Information about the various collateral types used by the DSS (as of Ethereum block 17121982).
@@
Standard ERC-20 Tokens ETH-A Ether 233580929 421276897 180% 15000000000 1.50% 145% 4.69%
@@
LP Tokens UNIV2USDCETH-A USDC/ETH 2139902 0 0% 1.50% 120% 0.04%
@@
Real-World Assets RWA001-A 6s Capital 14348036 15000000 105% 3.00% 0.29%
@@
PSMs PSM-USDC-A USD Coin 2118503205 2608700392 123% 10000000000 42.58%
@@
Appendix C Price Data Table
@@
Table 6. Data periods (times are in Greenwich Mean Time).
@@
Figure 6. The correlations between all 100 tokens in our dataset.
Appendix D Portfolio Compositions
@@
Table 7. DSS-based portfolio compositions.
@@
Table 8. Optimal portfolio compositions.
@@
Table 9. Optimal simulation portfolio if other stablecoins are included.